# The ML.ENERGY Benchmark: Toward Automated Inference Energy Measurement and Optimization

### v1

@@ Abstract
As the adoption of Generative AI in real-world services grow explosively, energy has emerged as a critical bottleneck resource. However, energy remains a metric that is often overlooked, under-explored, or poorly understood in the context of building ML systems. We present the ML.ENERGY Benchmark, a benchmark suite and tool for measuring inference energy consumption under realistic service environments, and the corresponding ML.ENERGY Leaderboard, which have served as a valuable resource for those hoping to understand and optimize the energy consumption of their generative AI services. In this paper, we explain four key design principles for benchmarking ML energy we have acquired over time, and then describe how they are implemented in the ML.ENERGY Benchmark. We then highlight results from the latest iteration of the benchmark, including energy measurements of 40 widely used model architectures across 6 different tasks, case studies of how ML design choices impact energy consumption, and how automated optimization recommendations can lead to significant (sometimes more than 40%) energy savings without changing what is being computed by the model. The ML.ENERGY Benchmark is open-source and can be easily extended to various customized models and application scenarios.
@@ 1 Introduction
Generative AI models have rapidly transitioned from research prototypes to real-world services such as ChatGPT [48], Character AI [4], Sora [49], and Midjourney [45]. However, exponential growth rarely continues without facing scaling bottlenecks; currently for generative AI, one of the most crucial bottlenecks is the energy bottleneck [13, 14, 43, 44, 46, 3]. That is, even with fleets of latest GPUs and exploding demand for ML compute, getting access to the energy necessary to power these systems is becoming increasingly costly, slow, and sometimes impossible. This particularly impacts serving real-world services as ML inference reportedly accounts for 80–90% of the total compute demand [27, 10, 51, 50]. Left unaddressed, the energy bottleneck will not only hinder AI research and development progress [26], but also lead to energy being either squeezed out of existing electricity grids and impacting availability and price [3], or generated from power sources like fossil fuels [29].
However, despite its growing importance, energy remains a secondary consideration compared to traditional optimization objectives like time and accuracy [57]. How much energy does a model consume during inference? What is the right way for energy measurement and accounting during execution, let alone optimization? To bridge this gap, we launched the ML.ENERGY Leaderboard, [^1] https://ml.energy/leaderboard the first inference energy leaderboard for modern generative AI models to the best of our knowledge. The Leaderboard has been gradually expanding in multiple dimensions to now include (1) 40 different generative AI model architectures across a wide range of tasks – including LLM [64] chat and coding, VLM [38] visual chat, and text-to-image, text-to-video, and image-to-video generation using Diffusion models [20, 12] – and (2) more up-to-date hardware and software stacks following rapid advancements in each area.
In this paper, we share the design principles we have established over time (Section 2) and present the ML.ENERGY Benchmark that embodies these principles (Section 3). It provides two key functionalities as a tool:
@@
Finally, we highlight notable results from the most recent iteration of the ML.ENERGY Leaderboard, shedding light on (1) how much energy consumption varies across different generative AI models and tasks, (2) the complex trade-offs that involve energy, time, and model architecture design, and (3) the energy savings opportunity unlocked by automated optimization (Section 4).
The ML.ENERGY Benchmark is open-source on GitHub, [^2] https://github.com/ml-energy/leaderboard and the ML.ENERGY Leaderboard allows everyone to browse full results from the ML.ENERGY Benchmark. The benchmark is designed to be easily extensible, allowing users to benchmark their models and compare them to others without the heavy lifting of building their own benchmarking dataset, runtime, and analysis infrastructure.
@@ 2.1 Generalizability and Portability
  • GPUs are the dominant worker and energy consumer in a system running ML services, accounting for 50–70% of the total provisioned power in the datacenter [50].
@@ 2.2 Representing Real-World Deployments
Benchmarking results often inform real-world deployment optimizations, are used to plan future energy usage, affect the design of new hardware and software systems, and serve as base numbers for long term projections that affect policymaking. Therefore, it is crucial that our measurements represent those from real-world deployments as closely as possible.
@@
  • During our measurement, we directly run or closely mimic the state of a well-utilized serving system during long term deployment. This allows us to capture the steady state energy consumption of the service while using a fixed-size benchmarking dataset.
@@ 2.4 Actionable Measurement Results
While energy measurements are useful in themselves, they are even more useful when they lead to actionable insights and recommendations. For instance, how much is the potential energy savings of your model without sacrificing accuracy or latency? If your service allows inference to run 10% slower, what is the energy-optimal configuration, and how much is the potential energy savings?
@@
The ML.ENERGY Benchmark allows users to provide computation latency constraints specific to their application scenario (e.g., LLM average Time Per Output Token), and will automatically recommend (1) the energy-optimal configuration that meets the latency constraints, and (2) the expected amount of energy savings. Due to the generalizability of our measurements (Section 2.1), these recommendations generalize to optimizing a wide range of systems.
@@ 3.2 Energy Measurement and Service-Aware Energy Accounting
Figure 2: LLM inference server and per-request energy accounting. The steady state is defined as the period when batch size is saturated at the server’s maximum configured batch size, and measurements during the steady state represent that of a well-utilized serving system during long-term deployment.
@@
Request-level energy accounting is less straightforward for LLM inference, because iteration-level batching [71] is an essential optimization in any realistic LLM serving system [34]. Figure 2 shows how requests are served by a serving system implementing iteration-level batching and how the ML.ENERGY Benchmark performs energy accounting. Because the beginning and end of each request are often not aligned with each other, finding each request’s individual energy consumption is non-trivial. For this, we first submit all requests in the request dataset, and as the system runs, identify the steady state as the time period where the batch size is saturated at the server’s maximum configured batch size. This steady state is designed to closely approximate the state of a serving system when it is well-utilized during long-term deployment. Particularly, when the system is ramping up initially with a full queue or ramping down at the end with an empty queue, the server runs with a smaller batch size and does not exhibit the same energy amortization benefits as the steady state. With this, we can derive the average per-request energy consumption with:
@@
In essence, we compute the average energy consumption per token during the steady state and multiply it by the average number of output tokens to derive the average per-request energy consumption.
@@ 3.3 Automated Optimization Recommendation
Our goal is to provide actionable insights beyond just energy measurements (Section 2.4) by recommending energy-optimal configurations for a given model and task. Central to the optimization recommendation is the construction of the Pareto frontier of energy vs. time, which is a collection of configurations where there is no other configuration that leads to both lower energy and lower time. Then, the energy-optimal configuration is selected based on user-specified latency constraints.
Latency constraints inherently depend on the user’s or application’s needs. For example, for image generation with Diffusion models, computation results are useful only when the full image is generated, so latency constraints would be specified in terms of the time to generate the whole image. On the other hand, for LLM text generation for chat, output tokens are streamed to users (either in written text or synthesized speech) as they are generated, so as long as the average time per output token is at least as fast as the users’ reading or listening speed, user experience will not be affected [39]. However, for LLM text generation for coding, where code is likely only useful when it is fully generated, latency constraints would be specified in terms of the time to generate the whole snippet, similar to the case of image generation. Given the latency constraints, the time–energy Pareto frontier is used to suggest the minimum-energy configuration that satisfies the latency constraint.
@@ 3.4 Extending the Benchmark
The ML.ENERGY Benchmark already supports various popular architectures like Llama [64], LLaVA [38], Stable Diffusion [20], and Stable Video Diffusion [12] (See Appendix A for a full list). Models that are fine-tuned based on already-supported models work as is. Models with different architectures are also supported as is as long as they are supported by the underlying runtime, like vLLM [34], which supports arbitrary LLMs provided by Hugging Face Transformers.
@@
For each task (e.g., LLM text generation for chat), the ML.ENERGY Benchmark provides a default request dataset that contains a set of inputs representative of real-world usage (See Appendix A for a full list). Users can also provide their own request dataset, which is then used to invoke the runtime and measure energy consumption.
@@
Energy consumption (Joules) reported by the ML.ENERGY Benchmark is a fundamental quantity that can be used to derive other useful metrics.
  • Average power draw (Watts): Average power draw can be calculated by dividing total energy consumption by the time taken to run the benchmark. This is useful for power provisioning.
  • Monetary cost (\$): Monetary cost can be calculated by multiplying energy consumption by the cost of electricity in the region and time frame in which the benchmark was run.
  • Operational carbon emissions (gCO2e): This quantity estimates the greenhouse gas emissions associated with the electricity consumed. It can be calculated by multiplying energy consumption by the carbon intensity of the particular region and time frame in which the benchmark was run.
@@ 4 Results Highlight
In this section, we highlight notable results from the ML.ENERGY Benchmark; the full set of results is available on the ML.ENERGY Leaderboard. [^3] https://ml.energy/leaderboard The most recent iteration of the benchmark and leaderboard presents energy measurements across 40 models and 6 tasks (See Appendix A for a full list). We ran the benchmark on NVIDIA A100 (40 GB) and H100 (80 GB) GPUs, each using AWS p4d.24xlarge and p5.48xlarge instances, respectively, and used vLLM [34] and Diffusers [68] as the inference runtime. In the following, we first present energy measurement results and discuss implications (Section 4.1), and then provide deeper understanding by showing how model architecture choices affect their energy consumption (Section 4.2). Then, we present the energy savings opportunities from our automated optimization recommendations (Section 4.3).
@@ 4.1 Energy Measurements
The dimmed bars behind each solid bar in Figure 3 show the estimated energy consumption based on the GPU’s Thermal Design Power (TDP) instead of measuring the real GPU power consumption, which is a common practice [35, 65, 23, 6, 7, 42]. Estimations using TDP are nearly always an overestimation since it is rare for a GPU – or any computing device – to draw its maximum power at every moment in time. In fact, such an estimation can lead to a worst-case overestimation of energy consumption by a factor of 4.1 (CodeGemma 2B on H100 GPUs). Inaccuracies may be overlooked when they influence downstream decisions and projections, leading to misleading conclusions. Therefore, it is crucial to aim for more accurate measurements.
@@ 4.2 Energy Implications of ML Design Decisions
ML decisions reflected in model architectures and trained models impact energy consumption. For the interest of space, we defer system parameter implications on energy consumption to Appendix B.
@@
In Figure 3, we can see that energy consumption varies even among LLMs of similar sizes. This is because different LLMs generate responses of different length even when given the same prompt. Such differences in verbosity can be non-trivial; for instance, Mistral Large’s responses were on average 36% longer than that of Mixtral 8$\times$7B. As the number of output tokens equals the number of forward passes through the model, longer responses leads to a proportional increase in energy consumption. As humans are known to prefer longer responses [74], this potentially introduces a trade-off between energy consumption and user satisfaction.
@@
Generally, models with more parameters consume more energy, but this is not always the case. Figure 4 highlights the case of Phi-3 Mini (3.8B) and Small (7B) [21]. Even though Small has nearly twice the parameters, the left plot shows that the larger Small model can consume less energy than Mini as batch size grows. This happens because Mini uses Multi-Head Attention (MA) [67], whereas Small uses Grouped Query Attention (GQA) [8]. Due to this, Mini’s KV cache uses 3$\times$ more memory than Small, which prevents it from scaling to larger batch sizes and amortizing energy consumption across more generations.
@@
Figure 6 shows the energy consumption of Stable Diffusion XL (SDXL) [52] and SDXL Turbo [5]. On the left, while SDXL and SDXL Turbo have identical model sizes and architectures, their energy consumption is significantly different. This is because SDXL Turbo is tuned to generate smaller resolution images (512$\times$512) than SDXL (1024$\times$1024), which leads to different latent sizes and amounts of computation. On the right, it can be seen that the number of denoising steps linearly increases energy consumption, as one denoising step requires one forward pass through the model. While simple in isolation, these inference-time parameters lead to non-trivial design tradeoffs at the application-level. For instance, increasing the number of denoising steps may improve final image quality, but beyond some point, it may be virtually indistinguishable to human users. Also, generating images in lower resolution and then upscaling them with a separate super-resolution model (e.g., DAT [16]) may consume less energy end-to-end.

### v2

@@ Abstract
As the adoption of Generative AI in real-world services grow explosively, energy has emerged as a critical bottleneck resource. However, energy remains a metric that is often overlooked, under-explored, or poorly understood in the context of building ML systems. We present the ML.ENERGY Benchmark, a benchmark suite and tool for measuring inference energy consumption under realistic service environments, and the corresponding ML.ENERGY Leaderboard, which have served as a valuable resource for those hoping to understand and optimize the energy consumption of their generative AI services. In this paper, we explain four key design principles for benchmarking ML energy we have acquired over time, and then describe how they are implemented in the ML.ENERGY Benchmark. We then highlight results from the early 2025 iteration of the benchmark, including energy measurements of 40 widely used model architectures across 6 different tasks, case studies of how ML design choices impact energy consumption, and how automated optimization recommendations can lead to significant (sometimes more than 40%) energy savings without changing what is being computed by the model. The ML.ENERGY Benchmark is open-source and can be easily extended to various customized models and application scenarios.
@@ 1 Introduction
Generative AI models have rapidly transitioned from research prototypes to real-world services such as ChatGPT [56], Character AI [6], Sora [57], and Midjourney [50]. However, exponential growth rarely continues without facing scaling bottlenecks; currently for generative AI, one of the most crucial bottlenecks is the energy bottleneck [15, 16, 17, 48, 49, 51, 4, 38]. That is, even with fleets of latest GPUs and exploding demand for ML compute, getting access to the energy necessary to power these systems is becoming increasingly costly, slow, and sometimes impossible. This particularly impacts serving real-world services as ML inference reportedly accounts for 80–90% of the total compute demand [32, 12, 60, 58]. Left unaddressed, the energy bottleneck will not only hinder AI research and development progress [31], but also lead to energy being squeezed out of existing electricity grids and impacting availability and price [4].
However, despite its growing importance, energy remains a secondary consideration compared to traditional optimization objectives like time and accuracy. How much energy does a model consume during inference? What is the right way for energy measurement and accounting during execution, let alone optimization? To bridge this gap, we launched the ML.ENERGY Leaderboard, [^1] https://ml.energy/leaderboard the first inference energy leaderboard for modern generative AI models to the best of our knowledge. We have been gradually expanding the Leaderboard in multiple dimensions to now include (1) 40 different generative AI model architectures across a wide range of tasks – including Large Language Model (LLM) chat and coding, Vision–Language Model (VLM) visual chat, and text-to-image, text-to-video, and image-to-video generation using Diffusion models – and (2) more up-to-date hardware and software stacks following rapid advancements in each area.
In this paper, we share the design principles we have established over time (Section 2) and present the ML.ENERGY Benchmark that embodies them (Section 3). It provides two key functionalities:
@@
Finally, we highlight notable results from the early 2025 iteration of the ML.ENERGY Leaderboard, shedding light on (1) how energy consumption varies across different generative AI models and tasks, (2) the complex trade-offs that involve energy, time, and model architecture design, and (3) the energy savings opportunity unlocked by automated optimization (Section 4).
This paper describes the state of the ML.ENERGY Benchmark and Leaderboard as of early 2025. The latest version of the ML.ENERGY Benchmark is open-source on GitHub, [^2] https://github.com/ml-energy/benchmark and the ML.ENERGY Leaderboard allows everyone to browse full results from the latest ML.ENERGY Benchmark.
@@ 2.1 Generalizability and Portability
  • GPUs are the dominant worker and energy consumer in a system running ML services, accounting for 50–70% of the total provisioned power in the datacenter [58, 52, 53, 54].
@@ 2.2 Representing Real-World Deployments
Benchmarking results often inform real-world deployment optimizations, are used to plan future power capacity and energy usage, affect the design of new hardware and software systems, and serve as base numbers for long term projections that affect policymaking. Therefore, it is crucial that our measurements represent those from real-world deployments as closely as possible.
@@
  • During our measurement, we directly run or closely mimic the state of a serving system during long term deployment. This allows us to capture the steady state energy consumption of the service while using a fixed-size benchmarking dataset.
@@ 2.4 Actionable Measurement Results
While energy measurements are useful in themselves, they are even more useful when they lead to actionable insights and recommendations. For instance, how much is the potential energy savings of your model without sacrificing accuracy or latency? If your service intends to guarantee a specific generation latency deadline (e.g., 50 ms), what is the energy-optimal configuration, and how much is the potential energy savings?
@@
The ML.ENERGY Benchmark allows users to provide computation latency constraints specific to their application scenario (e.g., LLM average Time Per Output Token), and will automatically recommend (1) the energy-optimal configuration that meets the latency constraints, and (2) the expected amount of energy savings. Due to the generalizability of our measurements (Section 2.1), these recommendations inform the optimization of a wide range of systems.
@@ 3.2 Energy Measurement and Service-Aware Energy Accounting
Figure 2: LLM inference server and per-request energy accounting. The steady state is defined as the period when batch size is saturated at the server’s maximum configured batch size, and measurements during the steady state represent that of a serving system during long-term deployment.
@@
Request-level energy accounting is less straightforward for LLM inference, because iteration-level batching [82] is an essential optimization in any realistic, production-grade LLM serving system [39]. Figure 2 shows how requests are served by a serving system implementing iteration-level batching and how the ML.ENERGY Benchmark performs energy accounting. Because the beginning and end of each request are often not aligned with each other, finding each request’s individual energy consumption is non-trivial. For this, we first submit all requests in the request dataset, and as the system runs, identify the steady state as the time period where the batch size is saturated at the server’s maximum configured batch size. This steady state is designed to closely approximate the state of a serving system when it is well-utilized during long-term deployment. Particularly, when the system is ramping up initially with a full queue or ramping down at the end with an empty queue, the server runs with a smaller batch size and does not exhibit the same energy amortization benefits as the steady state. With this, we can derive the average per-request energy consumption with:
@@
In essence, we compute the average energy consumption per token during the steady state and multiply it by the average number of output tokens to derive the average per-request energy consumption. Individual requests’ energy consumption can also be computed by multiplying the average energy per token during the steady state by the number of output tokens for each request.
As we will see in Section 4, batch size is a critical configuration that significantly affects both generation time and energy consumption. By sweeping the batch size configuration, the ML.ENERGY benchmark can capture varying levels of system utilization and collect various operation points with different time and energy consumption.
@@ 3.3 Automated Optimization Recommendation
Our goal is to provide actionable insights beyond just energy measurements (Section 2.4) by recommending energy-optimal configurations for a given model and task. Central to the optimization recommendation is the construction of the Pareto frontier of energy vs. time, which is a collection of configurations where there are no other configurations that lead to both lower energy and lower time. Then, the energy-optimal configuration is selected based on user-specified latency constraints.
Latency constraints inherently depend on the user’s or application’s needs. For example, for image generation with Diffusion models, computation results are useful only when the full image is generated, so latency constraints would be specified in terms of the time to generate the whole image. On the other hand, for LLM text generation for chat, output tokens are streamed to users (either in written text or synthesized speech) as they are generated. As such, for user-facing conversational AI services, as long as the average time per output token is at least as fast as the users’ reading or listening speed, user experience will not be affected [44]. However, for LLM text generation for coding, where code is likely only useful when it is fully generated, latency constraints would be specified in terms of the time to generate the whole snippet, similar to the case of image generation. Given the latency constraints, the time–energy Pareto frontier is used to suggest the minimum-energy configuration that satisfies the latency constraint.
@@ 3.4 Extending the Benchmark
The ML.ENERGY Benchmark already supports various popular architectures like Llama [73], LLaVA [43], Stable Diffusion [25], and Stable Video Diffusion [14] (See Appendix A for a full list). Models that are fine-tuned based on already-supported models work as is. Models with different architectures should also work as is as long as they are supported by the underlying runtime, like vLLM [39], which supports arbitrary LLMs provided by Hugging Face Transformers.
@@
For each task (e.g., LLM text generation for chat), the ML.ENERGY Benchmark provides a default request dataset that contains a set of inputs representative of real-world usage (See Appendix A for a full list). Users can also provide their own request dataset, which can be used to invoke the runtime and measure energy consumption.
@@
Hardware.
As long as the runtime used by the ML.ENERGY Benchmark (e.g., vLLM) is capable of running on the target hardware and Zeus [2] can measure energy consumption on the target hardware (e.g., NVIDIA/AMD GPUs, Intel/AMD CPUs, Apple Silicon, NVIDIA Jetson platforms), the ML.ENERGY Benchmark can run on the target hardware as is.
@@
Energy is a fundamental physical quantity that can be used to derive other useful metrics, though these derived metrics are not automatically computed by default as they require context-specific information. Below, we describe how these metrics might be computed based on the benchmark’s outputs.
  • Average power draw (Watts): Average power draw over the steady state can be calculated by dividing total energy consumption during the steady state by the duration of the steady state.
  • Throughput per Watt: Work throughput, e.g., request or token generation throughput, divided by average power draw can describe how much service capacity can be extracted from the system given a power budget, which is a critical quantity for datacenter power planning [38].
  • Monetary cost (\$): The electricity cost of compute can be calculated by integrating over time the multiplication of energy consumption and the electricity price in the region and time instance. If there is a specific region and time frame the service is expected to run, choosing that electricity price can simulate the operational electricity cost of deployment. Electricity prices can be obtained from sources like OpenEI. [^3] https://openei.org/wiki/Utility_Rate_Database Calculating the electricity cost from energy is supported by Zeus [2], the measurement library of choice for the benchmark.
  • Operational carbon emissions (gCO2e): This quantity estimates the greenhouse gas emissions associated with the electricity consumed. It can be calculated by multiplying energy consumption by the carbon intensity (gCO2e/kWh) of the particular region and time frame in which the benchmark was run. Carbon intensity data can be obtained from sources like ElectricityMaps. [^4] https://electricitymaps.com/ This is also supported by Zeus [2], the energy measurement library employed by the benchmark.
@@ 4 Results Highlight
In this section, we highlight notable results from the ML.ENERGY Benchmark; the full set of results is available on the ML.ENERGY Leaderboard. [^5] https://ml.energy/leaderboard The early 2025 iteration of the benchmark and leaderboard presents energy measurements across 40 models and 6 tasks (See Appendix A for a full list). We ran the benchmark on NVIDIA A100 (40 GB) and H100 (80 GB) GPUs, each using AWS p4d.24xlarge and p5.48xlarge instances, respectively, and used vLLM [39] and Diffusers [77] as the inference runtime. In the following, we first present energy measurement results and discuss implications (Section 4.1), and then provide deeper understanding by showing how model architecture choices affect their energy consumption (Section 4.2). Then, we present the energy savings opportunities from our automated optimization recommendations (Section 4.3).
@@ 4.1 Energy Measurements
The dimmed bars behind each solid bar in Figure 3 show the estimated energy consumption based on the GPU’s Thermal Design Power (TDP) instead of measuring the real GPU power consumption, which is a common practice [40, 74, 28, 8, 9, 47]. Estimations using TDP are nearly always an overestimation since it is rare for a GPU – or any computing device – to draw its maximum power at every moment in time. In fact, such an estimation can lead to a worst-case overestimation of energy consumption by a factor of 4.1 (CodeGemma 2B on H100 GPUs). Inaccuracies may be overlooked when they influence downstream decisions and projections, leading to misleading conclusions. Accurate measurements that reflect production environments are crucial.
@@ 4.2 Energy Implications of ML Design Decisions
Model TP Max batch size
4 8 16 32 64
DeepSeek distilled Qwen 3 8B [23, 80] 1 9713.7 6010.1 4314.9 3340.8 2770.8
Phi 4 reasoning plus 15B [3] 1 19974.4 12389.6 9347.3 7634.9 7595.4
Qwen 3 32B [80] 2 26419.7 15168.3 9140.5 6165.5 4520.6
Qwen 3 235B-A22B thinking [80] 8 122523.1 86491.5 56720.4 40275.5 33096.4
Table 1: Energy per generation of reasoning models on GPQA [64] and NVIDIA H100 GPUs. TP is the tensor parallelism degree, which is also equal to the number of GPUs used.
@@
ML decisions reflected in model architectures and trained models impact energy consumption. For the interest of space, we defer systems implications on energy consumption to Appendix B.
@@
[…]
This is even more pronounced for reasoning models, which produce significantly more output tokens. Table 1 shows energy measurements for reasoning models on the GPQA dataset. Reasoning models produce one to two orders of magnitude more output tokens per request compared to standard chat models, significantly increasing energy consumption per generation. Additionally, due to their long output lengths, servers cannot run as large a batch size, preventing them from amortizing energy across more requests. This leads to higher energy per token as well, further increasing energy consumption. As long horizon reasoning and task decomposition become more common in real-world LLM-based applications, we expect this trend to continue.
@@
[…]
@@
Figure 6 shows the energy consumption of Stable Diffusion XL (SDXL) [61] and SDXL Turbo [7]. On the left, while SDXL and SDXL Turbo have identical model sizes and architectures, their energy consumption is significantly different. This is because SDXL Turbo is tuned to generate smaller resolution images ($\times512$) than SDXL ($\times1024$), which leads to different latent sizes and amounts of computation. On the right, it can be seen that the number of denoising steps linearly increases energy consumption, as one denoising step requires one forward pass through the model. While simple in isolation, these inference-time parameters lead to non-trivial design tradeoffs at the application-level. For instance, increasing the number of denoising steps may improve final image quality, but beyond some point, it may be virtually indistinguishable to human users. Also, generating images in lower resolution and then upscaling them with a separate super-resolution model (e.g., DAT [19]) may consume less energy end-to-end.